‏Assuming the tost of direct materials used is $ 1,500,000 , compute the total manufacturing costs using the information below . Raw materials inventory , January 1 $ 30,000 Raw materials inventory , December 31 60,000 Work in process , January 1 27,000 Work in process , December 31 18,000 Finished goods , January 1 60,000 Finished goods , December 31 48,000 1,500,000 Raw materials purchases 690,000 Direct labor 225,000 75,000 Factory utilities Indirect labor 500,000 630,000 Factory depreciation Operating expenses


1
Expert's answer
2021-05-03T10:51:28-0400

Cost of production produced=30 000+1500 000 -60 000+690 000+225 000+75 000+500 000+630 000=3 590 000

Total production cost=3 590 00+(18 000-27 000)=3 581 000

Commercial products=3 581 000+(48 000-60 000)=3 569 000